OpenStudy (anonymous):

a trigonometry question#5 given tan x/2=y find cos x/(1+sinx) in y term

OpenStudy (mertsj):

tan(x/2)= sinx/(1+cosx)

OpenStudy (mertsj):

Multiply numerator and denominator by (1-cosx)

OpenStudy (mertsj):

That will lead you to a good place.

OpenStudy (anonymous):

i'm not doing in that way. can you try and give me an anz

OpenStudy (anonymous):

where is "tan(x/2)= sinx/(1+cosx)" from?

OpenStudy (mertsj):

It's a half angle formula.

myininaya (myininaya):

i got another way

myininaya (myininaya):

just a sec

OpenStudy (anonymous):

i think it can be written in several form

OpenStudy (anonymous):

let me give you guy choices

myininaya (myininaya):

scanning...

OpenStudy (anonymous):

a.\[1-y^2/(1+y)^2\] b.\[1-y^2/1+y^2\] c.\[y^2-1/y^2+1\] d. \[y^2-1/(y+1)^2\]

myininaya (myininaya):

so we have \[\frac{\frac{1-y^2}{y^2+1}}{1+\frac{2y}{y^2+1}} \cdot \frac{y^2+1}{y^2+1}=\frac{1-y^2}{y^2+1+2y}=\frac{1-y^2}{(y+1)^2}\]

myininaya (myininaya):

after simplification

myininaya (myininaya):

I really couldn't get it the way mert was doing i found this way perfect though

myininaya (myininaya):

can you real my pdf?

myininaya (myininaya):

read*

OpenStudy (anonymous):

yes same here thx